A solution of hydrofluoric acid has a molar concentration of 0.01 mol/L, what is the pOH

of the solution? Ka=6.6x 10^-39


1
Expert's answer
2021-06-07T03:23:38-0400

According to the equation:

HF + H2O = H3O+ + F-

Ka = [H3O+] [F-] / [HF]

From here:

[H3O+] [F-] = [HF] × Ka

[H3O+]2 = [HF] × Ka = 0.01 mol/L × 6.6 × 10-4 mol/L = 6.6 × 10-6

[H3O+] = (6.6 × 10-6)1/2 = 8.1 × 10-3 mol/L

pH = -log[H3O+] = -log(8.1 × 10-3) = 2.1

pOH = 14 - pH = 14 - 2.1 = 11.9


Answer: 11.9